OpenStudy (anonymous):

y=arctan(x-sqrt(1+x^2))

OpenStudy (anonymous):

what u want done?

OpenStudy (anonymous):

\[\frac{ 1 }{ 1+(x-\sqrt{1+x^2)} }(1-\frac{ 1 }{ 2 }(1+x^2)^{-1/2})(2x)\]

OpenStudy (anonymous):

oh differentiate

OpenStudy (anonymous):

am I on the right track?

OpenStudy (anonymous):

oh (x-sqrt(1+x^2))^2

OpenStudy (anonymous):

that first term is squared on the bottom

OpenStudy (anonymous):

yep... you're on the right track...

OpenStudy (anonymous):

ok great, i will simplify

OpenStudy (anonymous):

\(\Large [arctan(f(x))]'=\frac{f'(x)}{1+[f(x)]^2} \)